# Small Gradient Norm Regret for Online Convex Optimization

### v2

@@ Abstract
This paper introduces a new problem-dependent regret measure for online convex optimization with smooth losses. The notion, which we call the $G^{\star}$ regret, depends on the cumulative squared gradient norm evaluated at the decision in hindsight $\sum_{t=1}^{T}\|\nabla\ell(x^{\star})\|^{2}$. We show that the $G^{\star}$ regret strictly refines the existing $L^{\star}$ (small loss) regret, and that it can be arbitrarily sharper when the losses have vanishing curvature around the hindsight decision. We establish upper and lower bounds on the $G^{\star}$ regret and extend our results to dynamic regret and bandit settings. As a byproduct, we refine the existing convergence analysis of stochastic optimization algorithms in the interpolation regime. Some experiments validate our theoretical findings.
@@ 1 Introduction
and in particular, the best decision in hindsight $x^{\star}\in\operatornamewithlimits{arg\,min}_{x\in\mathcal{X}}\sum_{t=1}^{T}\ell_{t}(x)$. Despite the powerful adversary, existing OCO algorithms can achieve sublinear regret under different assumptions on the losses $\ell_{t}$. For example, when the losses are convex and Lipschitz, [42] shows that online gradient descent (OGD) achieves $\mathcal{O}(\sqrt{T})$ regret; if $\ell_{t}$ are additionally strongly convex or exponentially concave, then $\mathcal{O}(\log T)$ is attainable using OGD or online Newton step [16]. These regret bounds are minimax optimal in the corresponding loss function classes [1].
Beyond the assumptions discussed above, another popular condition on the loss functions is smoothness, which is often satisfied by loss functions in machine learning tasks [30]. When the loss functions are non-negative and have Lipschitz continuous gradients, it is possible to obtain a problem-dependent regret $\mathcal{O}(\sqrt{L_{T}^{\star}})$ [30], where $L_{T}^{\star}=\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\ell_{t}(x)=\sum_{t=1}^{T}\ell_{t}(x^{\star})$ denotes the cumulative loss under the best decision in hindsight. This regret, often known as the $L^{\star}$-regret or small-loss bound [26, 39, 35, 38], can be significantly sharper than the standard $\mathcal{O}(\sqrt{T})$ regret upper bound if $L_{T}^{\star}=\mathcal{O}(1)$, while still maintaining the worst-case $\mathcal{O}(\sqrt{T})$ regret. Recent work has developed parameter-free online learning algorithms that achieve $\mathcal{O}(\sqrt{L_{T}^{\star}})$ regret [39, 40] as well as guarantees in terms of other problem-dependent quantities.
@@
  • We prove that existing algorithms that adapt to the $L^{\star}$ regret also adapt to the $G^{\star}$ regret, achieving an $\mathcal{O}(\sqrt{G_{T}^{\star}})$ bound with a matching lower bound. We further extend our analysis to dynamic regret and bandit settings, and showcase the implications of $G^{\star}$ regret in the context of stochastic optimization. We also show how to design smoothed versions of the loss functions to allow adaptation to the constrained small-loss bound.
@@ 1.1 Related literature
The $L^{\star}$ regret, also known as the small loss bound, can provide a tighter regret upper bound when the decision in hindsight achieves a small cumulative loss.
@@
where $x^{\star}\in\operatornamewithlimits{arg\,min}_{x\in\mathcal{X}}\sum_{t=1}^{T}\ell_{t}(x)$. In the literature, it is often assumed that $\{\ell_{t}\}_{t=1}^{T}$ are non-negative, and an upper bound $\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\ell_{t}(x)$, is typically used. For ease of later comparison, this paper adopts the definition (1), the recentered version of the $L^{\star}$ regret. Existing OCO algorithms can achieve $\mathcal{O}(\sqrt{L^{\star}_{T}})$ regret for general smooth convex functions and $\mathcal{O}(\log L^{\star}_{T})$ regret when the loss is exponentially concave or strongly convex [35, 30]. Recently, the $L^{\star}$ regret has also been extended to the adaptive regret setting [37, 38].
@@ 1.2 Notations
Throughout the paper, we use $\|\cdot\|$ to denote the Euclidean norm and $\langle\cdot,\cdot\rangle$ to denote the Euclidean inner product. Given a closed convex set $\mathcal{X}$, $\Pi_{\mathcal{X}}[\cdot]$ denotes the orthogonal projection onto $\mathcal{X}$. The diameter of a set is defined by $\mathrm{diam}(\mathcal{X})\coloneqq\sup_{x,y}\|x-y\|$. We use $\mathbb{S}_{n}\coloneqq\{x\in\mathbb{R}^{n}:\|x\|=1\}$ to denote the unit sphere centered at the origin. For brevity, we use $\inf\ell\coloneqq\inf_{x\in\mathbb{R}^{n}}\ell(x)$ to denote the infimum of $\ell$ over $\mathbb{R}^{n}$. We define $n]\coloneqq\{1,\ldots,n\}$.
@@ 2 Small gradient norm regret in online convex optimization
This section defines the $G^{\star}$ regret and compares it with the existing problem-dependent regret bounds. We also show that several existing OCO algorithms achieve $\mathcal{O}(\sqrt{G^{\star}})$ regret, accompanied by a matching lower bound. The following assumptions will be frequently invoked throughout the paper.
@@ 2. A2:
A1 looks slightly different from, but is equivalent to, the standard smoothness assumption $\ell_{t}(x)-\ell_{t}(y)-\langle\nabla\ell_{t}(y),x-y\rangle\leq\frac{L}{2}\|x-y\|^{2}$: an $L$-smooth convex function on $\mathbb{R}^{n}$ is $\frac{1}{L}$-strongly convex in the dual space. It yields a tighter lower bound than that provided by convexity alone [25]. This sharper lower bound is central to the derivation of the $G^{\star}$ regret.
@@
The derivation of A1 from $L$-smoothness requires $\ell_{t}$ to be defined over the whole space. This assumption can be slightly relaxed to the case where $\mathcal{X}\subsetneq\text{int}\operatornamewithlimits{dom}\ell_{t}$. We discuss this extension in the appendix.
@@ 2.1 $G^{\star}$ regret bound
The $G^{\star}$ regret inherits the translation invariance of the gradient and does not require the losses to be non-negative or lower bounded. In particular, for online linear optimization with $\ell_{t}(x)=\langle g_{t},x\rangle$, we have
@@
It is feasible to define the $G^{\star}$ regret via the gradient-norm minimization problem $\min_{x\in\mathcal{X}}G_{T}(x)$. However, the minimizer of $G_{T}$ need not coincide with the point $x^{\star}$ that minimizes the cumulative losses $\sum_{t=1}^{T}\ell_{t}$, and the resulting regret bounds would not be directly comparable to $L_{T}^{\star}$. For ease of comparison, we define $G_{T}^{\star}$ as the $x^{\star}$ that minimizes cumulative losses. This choice suffices to demonstrate the claimed tightness results.
@@
The $G^{\star}$ regret measures the cumulative deviation from the unconstrained stationarity condition $\nabla\ell_{t}(x)=0$. It lower bounds the $L^{\star}$ (assuming it exists) regret by self-boundedness [30]:
@@
Under A1 and A2, we always have $G^{\star}_{T}\leq 2L(L^{\star}_{T})$. Moreover, there exist loss sequences
@@
Proposition˜2.1 shows that small $L^{\star}$ regret implies small $G^{\star}$ regret, but the converse may not hold. Moreover, the $G^{\star}$ (or $L^{\star}$) regret and the gradient variation bound cannot imply each other. The remainder of this paper focuses on the $L^{\star}$ and $G^{\star}$ regret. Below are examples where the $G^{\star}$ regret can be substantially smaller than $L^{\star}$.
@@
Consider logistic regression with cross-entropy loss $\ell_{t}(x)=\log(1+\exp(-y_{t}\langle a_{t},x\rangle))$, where $\|a_{t}\|=1$ is a data vector and $y_{t}\in\{1,-1\}$. Then $\nabla\ell_{t}(x)=\tfrac{-y_{t}a_{t}}{1+\exp(y_{t}\langle a_{t},x\rangle)}$ and $\|\nabla\ell_{t}(x)\|^{2}\leq\ell_{t}(x)^{2}=o(\ell_{t}(x))$ whenever $\ell_{t}(x)\rightarrow 0$. The observation generalizes to multi-class classification.
@@
Consider online linear regression with loss $\ell_{t}(x)=\frac{1}{p}|r_{t}(x)|^{p}$ for $p\geq 2$, where the residual $r_{t}(x)=\langle a_{t},x\rangle-b$. This loss is flat near its optimum, and the squared gradient norm vanishes faster than the function value: $\nabla\ell_{t}(x)=a_{t}|r_{t}(x)|^{p-2}r_{t}(x)$ and $\tfrac{\|\nabla\ell_{t}(x)\|^{2}}{\ell_{t}(x)}=p\|a_{t}\|^{2}|r_{t}(x)|^{p-2}=p^{\frac{2p-2}{p}}\|a_{t}\|^{2}\ell_{t}(x)^{\frac{p-2}{p}}$. Then $\|\nabla\ell_{t}(x)\|^{2}=o(\ell_{t}(x))$ whenever $r_{t}(x)\rightarrow 0$.
@@
The above examples share the feature that the loss function becomes increasingly flat in a neighborhood of the optimum. For such loss functions, the $G^{\star}$ regret is typically much tighter than the $L^{\star}$ regret in the (near-)interpolation regime where $x^{\star}$ is nearly optimal for each of the losses.
@@
The $L^{\star}$ regret can be obtained by directly assuming self-boundedness of the losses (5). Self-boundeness alone is weaker than $L$-smoothness and covers certain nonsmooth functions [26]. On the other hand, it is unclear whether (3) can be generalized to handle nonsmooth losses. We leave further investigations to future work.
@@ 2.2 Upper and lower bound for the $G^{\star}$ regret
where (6) uses non-expansiveness of projection, (7) applies convexity $\ell_{t}(x)\geq\ell_{t}(x^{t})+\langle\nabla\ell_{t}(x^{t}),x-x^{t}\rangle$, and the $\|\nabla\ell_{t}(x^{t})\|^{2}$ term is further upperbounded in (8) using self-boundeness of (non-negative) $L$-smooth functions. However, a smooth function over $\mathbb{R}^{n}$ (A1) satisfies a stronger inequality
@@
A re-arrangement gives the standard regret inequality, and the regret bound follows from telescoping.
@@
Under A1 and A2, online gradient descent $x^{t+1}=\Pi_{\mathcal{X}}[x^{t}-\eta\nabla\ell_{t}(x^{t})]$ with $\eta\in(0,\tfrac{1}{L})$ satisfies
@@
The learning rate range in Theorem˜2.1 can be slightly enlarged to $(0,\frac{2}{L})$ at the cost of re-introducing $L_{T}(x)$:
@@
Under the same assumptions as Theorem˜2.1, online gradient descent $x^{t+1}=\Pi_{\mathcal{X}}[x^{t}-\eta\nabla\ell_{t}(x^{t})]$ with $\eta\in(0,\tfrac{2}{L})$ satisfies $\rho_{T}(x)\leq\tfrac{\|x^{1}-x\|^{2}}{(2-\eta L)\eta}+\tfrac{\eta}{2-\eta L}[LL_{T}(x)+\tfrac{1}{2-\eta L}G_{T}(x)]$ for any $x\in\mathcal{X}$.
Theorem˜2.1 resembles the guarantees of the $L^{\star}$ regret [26]. Using $G^{\star}_{T}\leq 2LL^{\star}_{T}$ from Proposition˜2.1, we immediately obtain guarantees in terms of the $L^{\star}$ regret: $\sum_{t=1}^{T}\ell_{t}(x^{t})-\ell_{t}(x)\leq\max\{2LD^{2},\sqrt{2LL^{\star}_{T}}D\}$. Besides, it always holds that $\sqrt{G^{\star}_{T}}=\mathcal{O}(\sqrt{T})$, retaining a worst-case $\mathcal{O}(\sqrt{T})$ regret.
@@
Since the conversion applies to any online algorithm with similar regret guarantees on the linearized losses, we can relax the bounded domain assumption using the recent results from unconstrained online learning [5] and obtain a $\rho_{T}(x)\leq\mathcal{O}(\|x\|\sqrt{G_{T}(x)}+\max_{t\in[T]}\|\nabla\ell_{t}(x^{t})\|^{2}+(L+1)\|x\|^{2})$ regret result.
@@
Finally, since $G^{\star}_{T}$ is equivalent to cumulative squared gradient norm for linear losses, the $\Omega(\sqrt{T})$ lower bound for online linear optimization implies an $\Omega(\sqrt{G^{\star}_{T}})$ regret lower bound.

### v3

@@ Abstract
This paper introduces a new problem-dependent regret measure for online convex optimization with smooth losses. The notion, which we call the $G^{\star}$ regret, depends on the cumulative squared gradient norm evaluated at the decision in hindsight. We show that the $G^{\star}$ regret strictly refines the existing $L^{\star}$ (small loss) regret, and that it can be arbitrarily sharper when the losses have vanishing curvature around the hindsight decision. We establish upper and lower bounds on the $G^{\star}$ regret and extend our results to dynamic regret and bandit settings. As a byproduct, we refine the existing convergence analysis of stochastic optimization algorithms in the interpolation regime. Some experiments validate our theoretical findings.
@@ 1 Introduction
and in particular, the best decision in hindsight $x^{\star}\in\operatornamewithlimits{arg\,min}_{x\in\mathcal{X}}\sum_{t=1}^{T}\ell_{t}(x)$. Despite the powerful adversary, existing OCO algorithms can achieve sublinear regret under different assumptions on the losses $\ell_{t}$. For example, when the losses are convex and Lipschitz continuous, [41] shows that online gradient descent (OGD) achieves $\mathcal{O}(\sqrt{T})$ regret; if $\ell_{t}$ are additionally strongly convex or exponentially concave, then $\mathcal{O}(\log T)$ regret is attainable using OGD or online Newton step [16]. These regret bounds are minimax optimal in terms of the corresponding loss function classes [1].
Beyond the assumptions discussed above, another popular condition on the loss functions is smoothness, which is often satisfied by loss functions in machine learning tasks [29]. When the loss functions are non-negative and have Lipschitz continuous gradients, it is possible to obtain a problem-dependent regret $\mathcal{O}(\sqrt{L_{T}^{\star}})$ [29], where $L_{T}^{\star}=\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\ell_{t}(x)=\sum_{t=1}^{T}\ell_{t}(x^{\star})$ denotes the cumulative loss under the best decision in hindsight. This regret, often known as the $L^{\star}$-regret or small loss bound [25, 38, 34, 37], can be significantly sharper than the standard $\mathcal{O}(\sqrt{T})$ regret upper bound if $L_{T}^{\star}=\mathcal{O}(1)$, while still maintaining the worst-case $\mathcal{O}(\sqrt{T})$ regret. Recent work has developed parameter-free online learning algorithms that achieve $\mathcal{O}(\sqrt{L_{T}^{\star}})$ regret [38, 39] as well as guarantees in terms of other problem-dependent quantities.
@@
  • We prove that existing algorithms that adapt to the $L^{\star}$ regret also adapt to the $G^{\star}$ regret, achieving an $\mathcal{O}(\sqrt{G_{T}^{\star}})$ bound with a matching lower bound. We further extend our analysis to dynamic regret and bandit settings, and showcase the implications of $G^{\star}$ regret in the context of stochastic optimization. We also show how to design smoothed versions of the loss functions to allow adaptation to the constrained small loss bound.
@@ 1.1 Related literature
The $L^{\star}$ regret, also known as the small loss bound, can provide a tighter regret upper bound when the decision in hindsight achieves a small cumulative loss
@@
where $x^{\star}\in\operatornamewithlimits{arg\,min}_{x\in\mathcal{X}}\sum_{t=1}^{T}\ell_{t}(x)$. In the literature, it is often assumed that $\{\ell_{t}\}_{t=1}^{T}$ are non-negative, and an upper bound $\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\ell_{t}(x)$, is typically used. For ease of later comparison, this paper adopts the definition (1), the recentered version of the $L^{\star}$ regret. Existing OCO algorithms can achieve $\mathcal{O}(\sqrt{L^{\star}_{T}})$ regret for general smooth convex functions and $\mathcal{O}(\log L^{\star}_{T})$ regret when the loss is exponentially concave or strongly convex [34, 29]. Recently, the $L^{\star}$ regret has also been extended to the adaptive regret [36, 37].
@@ 1.2 Notations
Throughout the paper, we use $\|\cdot\|$ to denote the Euclidean norm and $\langle\cdot,\cdot\rangle$ to denote the Euclidean inner product. Given a closed convex set $\mathcal{X}$, $\Pi_{\mathcal{X}}[\cdot]$ denotes the orthogonal projection onto $\mathcal{X}$. The diameter of a set is defined by $\mathrm{diam}(\mathcal{X})\coloneqq\max_{x,y\in\mathcal{X}}\|x-y\|$. We use $\mathbb{S}_{n}\coloneqq\{x\in\mathbb{R}^{n}:\|x\|=1\}$ to denote the unit sphere centered at the origin. For brevity, we use $\inf\ell\coloneqq\inf_{x\in\mathbb{R}^{n}}\ell(x)$ to denote the infimum of $\ell$ over $\mathbb{R}^{n}$. We define $n]\coloneqq\{1,\ldots,n\}$.
@@ 2 Small gradient norm regret in online convex optimization
This section defines the $G^{\star}$ regret and compares it with the existing problem-dependent regret bounds. We also show that several existing OCO algorithms achieve $\mathcal{O}(\sqrt{G^{\star}_{T}})$ regret, accompanied by a matching lower bound. The following assumptions will be frequently invoked.
@@ 2. A2:
The assumption A1 looks slightly different from, but is equivalent to, the standard smoothness assumption $\ell_{t}(x)-\ell_{t}(y)-\langle\nabla\ell_{t}(y),x-y\rangle\leq\frac{L}{2}\|x-y\|^{2}$: an $L$-smooth convex function on $\mathbb{R}^{n}$ is $\frac{1}{L}$-strongly convex in the dual space. It yields a tighter lower bound than that provided by convexity alone [24]. This sharper lower bound is central to the derivation of the $G^{\star}$ regret.
@@
Deriving A1 from $L$-smoothness requires $\ell_{t}$ to be defined over the whole space. This assumption can be slightly relaxed to the case where $\mathcal{X}\subsetneq\text{int}\operatornamewithlimits{dom}\ell_{t}$ (see Section˜A.1). Note that the traditional setting of $L^{\star}$ regret also requires the losses to be defined over $\mathbb{R}^{n}$.
@@ 2.1 $G^{\star}$ regret bound
The $G^{\star}$ regret inherits the translation invariance of the gradient and does not require the losses to be non-negative or lower bounded. In particular, for online linear optimization with $\ell_{t}(x)=\langle g_{t},x\rangle$,
@@
One may define the $G^{\star}$ regret via the gradient-norm minimization problem $\min_{x\in\mathcal{X}}G_{T}(x)$. However, the minimizer of $G_{T}$ need not coincide with the point $x^{\star}$ that minimizes the cumulative losses $\sum_{t=1}^{T}\ell_{t}$, and the resulting regret bounds would not be directly comparable to $L_{T}^{\star}$. For ease of comparison, we define $G_{T}^{\star}$ as the $x^{\star}$ that minimizes cumulative losses. This choice suffices to demonstrate the claimed tightness results.
@@
The $G^{\star}$ regret measures the cumulative deviation from the unconstrained stationarity condition $\nabla\ell_{t}(x)=0$. It lower bounds the $L^{\star}$ regret (assuming it exists) by self-boundedness [29]:
@@
Under A1 and A2, $G^{\star}_{T}\leq 2L(L^{\star}_{T})$. Moreover, there exist loss sequences
@@
Proposition˜2.1 shows that small $L^{\star}$ regret implies small $G^{\star}$ regret, but the converse may not hold. Moreover, the $G^{\star}$ (or $L^{\star}$) regret and the gradient variation bound cannot imply each other. The rest of this paper focuses on the $L^{\star}$ and $G^{\star}$ regret. Below are examples where the $G^{\star}$ regret can be substantially smaller than $L^{\star}$. They all satisfy $\inf\ell_{t}=0$, and we drop $\inf\ell_{t}$ for brevity.
@@
Consider logistic regression with cross-entropy loss $\ell_{t}(x)=\log(1+\exp(-y_{t}\langle a_{t},x\rangle))$, where $\|a_{t}\|=1$ is a data vector and $y_{t}\in\{1,-1\}$. Then
$\nabla\ell_{t}(x)=\tfrac{-y_{t}a_{t}}{1+\exp(y_{t}\langle a_{t},x\rangle)}\quad\text{and}\quad\|\nabla\ell_{t}(x)\|^{2}\leq\ell_{t}(x)^{2}=o(\ell_{t}(x))$
whenever $\ell_{t}(x)\rightarrow 0$. The observation generalizes to multi-class classification.
@@
Consider online linear regression with loss $\ell_{t}(x)=\tfrac{1}{p}|r_{t}(x)|^{p}$ for $p\geq 2$, where the residual $r_{t}(x)=\langle a_{t},x\rangle-b$. This loss is flat near the optimum, and the squared gradient norm vanishes faster than the function value: $\nabla\ell_{t}(x)=a_{t}|r_{t}(x)|^{p-2}r_{t}(x)$ and
$\tfrac{\|\nabla\ell_{t}(x)\|^{2}}{\ell_{t}(x)}=p\|a_{t}\|^{2}|r_{t}(x)|^{p-2}=p^{\frac{2p-2}{p}}\|a_{t}\|^{2}\ell_{t}(x)^{\frac{p-2}{p}}.$
Then $\|\nabla\ell_{t}(x)\|^{2}=o(\ell_{t}(x))$ whenever $r_{t}(x)\rightarrow 0$.
@@
The above examples share the feature that the loss functions are flat in a neighborhood of the optimum. For such loss functions, the $G^{\star}$ regret is typically much tighter than the $L^{\star}$ regret in the (near-)interpolation regime where $x^{\star}$ is nearly optimal for each of the losses.
@@
The $L^{\star}$ regret can be obtained by directly assuming self-boundedness of the losses (5). Self-boundedness alone is weaker than $L$-smoothness and covers certain nonsmooth functions [25]. On the other hand, it is unclear whether (3) can be generalized to handle nonsmooth losses. We leave further investigations to future work.
@@ 2.2 Upper and lower bound for the $G^{\star}$ regret
where (6) uses non-expansiveness of projection, (7) applies convexity $\ell_{t}(x)\geq\ell_{t}(x^{t})+\langle\nabla\ell_{t}(x^{t}),x-x^{t}\rangle$, and the $\|\nabla\ell_{t}(x^{t})\|^{2}$ term is further upperbounded in (8) using self-boundedness of (non-negative) $L$-smooth functions. However, a smooth function over $\mathbb{R}^{n}$ (A1) satisfies a stronger inequality
@@
A re-arrangement gives the standard regret inequality, and the regret bound follows by telescoping.
@@
Under A1 and A2, OGD: $x^{t+1}=\Pi_{\mathcal{X}}[x^{t}-\eta\nabla\ell_{t}(x^{t})]$ with $\eta\in(0,\tfrac{1}{L})$ satisfies
@@
The learning rate range in Theorem˜2.1 can be enlarged to $(0,\frac{2}{L})$ at the cost of re-introducing $L_{T}(x)$:
@@
Under A1 and A2, OGD: $x^{t+1}=\Pi_{\mathcal{X}}[x^{t}-\eta\nabla\ell_{t}(x^{t})]$ with $\eta\in(0,\tfrac{2}{L})$ satisfies $\rho_{T}(x)\leq\tfrac{\|x^{1}-x\|^{2}}{(2-\eta L)\eta}+\tfrac{\eta}{2-\eta L}[LL_{T}(x)+\tfrac{1}{2-\eta L}G_{T}(x)]$ for any $x\in\mathcal{X}$.
Theorem˜2.1 resembles the guarantees of the $L^{\star}$ regret [25]. Using $G^{\star}_{T}\leq 2LL^{\star}_{T}$ from Proposition˜2.1, we obtain guarantees of the $L^{\star}$ regret: $\sum_{t=1}^{T}\ell_{t}(x^{t})-\ell_{t}(x)\leq\max\{2LD^{2},\sqrt{2LL^{\star}_{T}}D\}$. Besides, it always holds that $\sqrt{G^{\star}_{T}}=\mathcal{O}(\sqrt{T})$, retaining a worst-case $\mathcal{O}(\sqrt{T})$ regret.
@@
Since the conversion applies to any online algorithm with similar regret guarantees on the linearized losses, we can relax the bounded domain assumption using the recent results from unconstrained online learning (for example, [6]) and obtain a $\rho_{T}(x)\leq\mathcal{O}(\|x\|\sqrt{G_{T}(x)}+\max_{t\in[T]}\|\nabla\ell_{t}(x^{t})\|^{2}+(L+1)\|x\|^{2})$ regret result.
@@
Since $G^{\star}_{T}$ is equivalent to cumulative squared gradient norm for linear losses, the $\Omega(\sqrt{T})$ lower bound for online linear optimization implies an $\Omega(\sqrt{G^{\star}_{T}})$ regret lower bound.